I need help with this problem. I am not sure if this problem falls under the category I put it in but I don't really know the category it is in:
1-2/3y=12/5-y/5

Answer by whisper(27)

To solve this we will use two processes. We will need to separate constants to one side of the equation and also find common denominators. I am assuming that is (2/3)y
1 - (2/3)y = 12/5 - y/5
1 - 12/5 -(2/3)y = 12/5 - 12/5 - y/5
5/5 - 12/5 - (2/3)y = 0 - y/5
-7/5 - (2/3)y = -y/5
-7/5 - (2/3)y + (2/3)y = -y/5 + (2/3)y
-7/5 + 0 = -y/5 + (2/3)y
-7/5 = -y/5 + (2/3)y
At this point I am going to "clear" the denominator by multiplying through by 5.
5(-7/5) = 5[-y/5 + (2/3)y]
-7 = -y + (10/3)y
Find a common denominator for the y coefficients:
-7 = -(3/3)y + (10/3)y
-7 =-(7/3)y
Multiply both sides by the recipricol of -(7/3) which is (-3/7)
-7(-3/7)=-(7/3)[-(3/7)]
3 = 1y
3=y
Please don't hesitate to ask any further questions!